nineteen coins consisting of quarters and half-dollars amount to 6.50. find the amount of coins of each kind.
Found 2 solutions by stanbon, robertb:
Answer by stanbon(75887)

nineteen coins consisting of quarters and half-dollars amount to 6.50. find the amount of coins of each kind.
--------------
Quantity Eq.:: q + h = 19
Value Eq:::: 25q+50h = 650 cents
-----------------------------------
Multiply thru 1st by 25 to get:
25q+25h = 25*19
---
Subtract that from the Value Eq and solve for "h":
25h = 650-25*19
25h = 175
h = 7 (# of half-dollar coins)
----
Solve for "q"
q+h = 19
q+7 = 19
q = 12 (# of quarters)
===========================
Cheers,
Stan H.
=============

Answer by robertb(5830)

0.25q + 0.50(19-q) = 6.50
==> 0.25q + 9.5 - 0.50q = 6.5
==> -0.25q = -3
==> q = 12, the number of quarters, and
19-q = 19-12 = 7, the number of half-dollars.